i need help finding x in the equation: 2(squareroot x) + 3 = x
Answer by mananth(16946)

i need help finding x in the equation: 2(squareroot x) + 3 = x
2sqrtx +3=x
2sqrtx=x-3
square both the sides of the equation
4x=(x-3)^2
4x=x^2-6x+9
x^2-4x-6x-9=0
x^2-10x-9=0
x^2-9x-x+9=0
x(x-9)-1(x-9)=0
(x-1)(x-9)=0